# A dynamical algorithm to compute hyperbolic Julia sets in polynomial time

### v1

@@ 3.2. Poly-time computability of $J$
Step 1(a). First, if $p$ is hyperbolic, we can build a collection of boxes on which an iterate of $p$ is expanding. That is, we invoke Lemma 3.1, which succeeds if $p$ is hyperbolic, to compute an integer $\nu$, dyadic rationals $\delta>0,L>1$, and the set $\mathcal{N}_{1}$ which is a collection of boxes on which $|Dp^{\nu}(z)|>L$. If $p$ is not hyperbolic, this step will never halt. If this step halts, then the algorithm has just proven that $p$ is hyperbolic (see Lemma 3.9 of [3]).
@@
Step 1(b). Next, we construct our “reference boxes” $\mathcal{N}_{2}$, such that $J$ is contained in the union of the boxes, each box contains a point of $J$, and $|Dg|>L>1$ on $\mathcal{N}_{2}$. We use the algorithm provided in Proposition 3.2. This proposition allows us to input a $b>0$ that is an upper bound for the side-length $\beta^{\prime}$ of the boxes in $\mathcal{N}_{2}$. We next provide the calculation defining the $b$ we need; the reason this value is needed will be demonstrated in the proof of Theorem 1.1.
First set $r=r(\delta):=3\delta/4.$ Now choose any $\gamma\leq 0.1$ (we see much later why that $\gamma$ is needed). Then use Equation 2.3 defining $\gamma_{r}(a)$ to determine $a$.
Now set $c=c(\gamma)=(1-3\gamma)(1-\gamma)/({1+\gamma})$ and set
(3.1) $b:=\min\left(\frac{\delta\ c(\gamma)}{16\sqrt{2}(1+\gamma)},a\right).$
Then invoke Proposition 3.2 and compute the collection $\mathcal{N}_{2}$ of boxes of computable side-length $0<\beta^{\prime}<b$. Recall also $\mathcal{N}(\mathcal{N}_{2},\delta)\subset\mathcal{N}_{1}$.
Step 1(c). In the following rationale for how to draw the $2^{N}$-picture of $J$, we need to require $N\geq N^{\prime}$ so that both $2^{-N-1}<a$ and $2^{-N}<\delta$. To draw a $2^{N}$-picture of $J$ for some $N<N^{\prime}$ we simply use the more precise $2^{N^{\prime}}$-approximation.
Consider any $N\geq N^{\prime}$. We perform the following for every ideal point $z^{\prime}$, a dyadic rational in $\mathbb{C}$ with coordinate denominators in $2^{N+2}$—except first, since $\mathcal{N}_{2}\supset J$, define $h_{J}(N,z^{\prime})=0$ for all $z^{\prime}\in\mathbb{C}\setminus[-R-1,R+1]^{2}$, where we calculated $R,R]^{2}$ a large box containing $J$ at the beginning of the algorithm, described in Lemma 3.1. Now, for each $z^{\prime}\in[-R-1,R+1]^{2}$, proceed as follows.
Remember $B(z^{\prime},r)$ denotes the ball (box in our norm) about $z^{\prime}$ of radius $r$. For ease of notation, set $\mathcal{U}^{\prime}:=\overline{B(z^{\prime},2^{-N-2})}$ and $2\mathcal{U}^{\prime}:=\overline{B(z^{\prime},2^{-N-1})}$.
@@
Step 2(a). We first calculate the maximum number of iterates that have to be considered. Define
@@
where $\lceil\cdot\rceil$ is the ceiling function, $\gamma>0$ and $c=c(\gamma)=(1-3\gamma)(1-\gamma)/({1+\gamma})$ are as in Step 1(b), $L$ is the lower bound on expansion of $|Dg|$ in $\mathcal{N}_{1}$, and $\beta^{\prime}$ is the side-length of the reference boxes in $\mathcal{N}_{2}$. It is important that $k_{N}$ is linear in $N$.
@@
We explained in Step 1(a) of Algorithm 3.3 (and in Section 3.1) that if Step 1(a) succeeds, then we have established that $p$ is hyperbolic, and built a collection of boxes $\mathcal{N}_{1}$ containing $J$ and on which $p$ is expanding, and a smaller collection of boxes $\mathcal{N}_{2}$ containing $J$ such that each box contains a point of $J$, and $\mathcal{N}_{2}\subset\mathcal{N}(\mathcal{N}_{1},\delta)$. (Note if $p$ is not hyperbolic, Step 1(a) never halts so we do not make it to the tests Step 1(d, e, or f) to report anything for any $z^{\prime}$, so this is why we do not need to assume $p$ is hyperbolic for the lemma, that is implied by the assumption the algorithm makes a report in Steps 1(d, e, or f).)
@@
Step 1(e) contained the second check. Since every box in $\mathcal{N}_{2}$ contains a point of $J$, if there is a reference box entirely contained in $2\mathcal{U}^{\prime}$, then there is a point of $J$ in $2\mathcal{U}^{\prime}=\overline{B(z^{\prime},2^{-N-1})},$ so $d(J,z^{\prime})\leq 2^{-N-1}$. So we are allowed to report $1$.
@@
(i): First we show $B^{e}(z^{\prime},r)\subset\mathcal{N}_{1}$, in particular $g$ is conformal on $B^{e}(z^{\prime},r)$, for $r=3\delta/4$. Since Step 1(d) did not triggered a report, we know $\mathcal{N}_{2}\cap\mathcal{U}^{\prime}\neq\emptyset$. Thus $d(z^{\prime},\mathcal{N}_{2})\leq 2^{-N-2}$. Recall that $\mathcal{N}_{1}\supset\mathcal{N}(\mathcal{N}_{2},\delta).$ Since $\delta>2^{-N^{\prime}}\geq 2^{-N},$ we have $2^{-N-2}<\delta/4,$ so $d(z^{\prime},\mathcal{N}_{2})\leq\delta/4,$ hence $B^{e}(z^{\prime},3\delta/4)\subset\mathcal{N}_{1}$. Finally, recall that $g$ is conformal on $\mathcal{N}_{1}$ since we have $|Dg(z)|>L>1>0$ there.
(ii): Next we show if $w^{\prime}\in\Omega_{L}$, then $d(w^{\prime},\Omega_{U})\geq\sigma_{N}$, and for any $i\in\mathbb{bN}$, $\frac{1}{2}{\rm diam}(g^{i}(B^{e}(w^{\prime},\sigma_{N}))\geq|Dg^{i}(z^{\prime})|_{e}\sigma_{N}c(\gamma).$ We aim at applying Proposition 2.12 and Lemma 2.13 for $g^{i}=(p^{\nu})^{i},r=(3\delta/4)$, $\gamma\leq 0.1$ and $a<r$ based on $\gamma_{r}(a)$ defined by Equation 2.3, $z_{0}=z^{\prime}$, $w_{0}=w^{\prime}$ any point of $\Omega_{L}$, and $\sigma=\sigma_{N}$ (see Figure 1, Right). Since we are in Step 2, we know $\mathcal{U}^{\prime}\cap\mathcal{N}_{2}\neq\emptyset$ so $z^{\prime}\in\mathcal{N}(\mathcal{N}_{2},\delta/2)\subset\mathcal{N}_{1}$ and $\Omega_{L}\subset\Omega_{U}\subset\mathcal{N}_{1}$ as well since $2^{-N-1}<\delta/2$. So since we chose $N^{\prime}$ so that $2^{-N^{\prime}-1}<a,$ we know $\Omega_{L}\subset B^{e}(z^{\prime},a)$. Hence $w^{\prime}\in B^{e}(z^{\prime},a)$, and Lemma 2.13 yields (ii).
@@
Observe if the difference of the radii between the underestimate of $g^{i}(\Omega_{U})$ and the overestimate of $g^{i}(\Omega_{L})$ is at least $\beta^{\prime}\sqrt{2},$ then we know the reference box about $g^{i}(w^{\prime})\in g^{i}(\Omega_{L})$ is contained entirely inside of the underestimate of $g^{i}(\Omega_{U})$. We show this is true. In fact, this difference in radii is at least $|Dg^{i}(z^{\prime})|_{e}4\sigma_{N}(1-\gamma)-|Dg^{i}(z^{\prime})|_{e}3\sigma_{N}(1+\gamma)=|Dg^{i}(z^{\prime})|_{e}\sigma_{N}(1-7\gamma).$ We know $c(\gamma)=(1-3\gamma)(1-\gamma)/(1+\gamma),$ so by an easy calculation, by choice of $\gamma\leq 0.1$, we have $(1-7\gamma)\geq c(\gamma)/2$. (Simplifying this inequality leads to a quadratic which opens down, with two roots, one negative, and the positive root has solution $\gamma=-(8-\sqrt{132}/34)\approx 0.1026213$, so any $\gamma\in(0,0.1]$ satisfies this desired inequality).
Thus, $|Dg^{i}(z^{\prime})|_{e}\sigma_{N}(1-7\gamma)\geq|Dg^{i}(z^{\prime})|_{e}\sigma_{N}c(\gamma)/2.$ So since by hypothesis this latter quantity is $\beta^{\prime}\sqrt{2}$, we know the the minimum difference between the underestimate of $g^{i}(\Omega_{U})$ to the boundary of the overestimate of $g^{i}(\Omega_{L})$, is more than the euclidean diameter of a reference box, and we conclude that the reference box about $g^{i}(w^{\prime})$ is contained entirely inside of the underestimate of $g^{i}(\Omega_{U})$.
@@
Now suppose $k\geq 2$ and we have moved through Steps 2(b),(c),(d), and (e) for $k$. Hence, the algorithm has made it through Steps for $1\leq j\leq k-1$ without halting. Thus:
@@
Suppose to the contrary that $g^{k-1}(z^{\prime})\notin\mathcal{N}(\mathcal{N}_{1},\delta/2)$. We know $z^{\prime}\in\mathcal{N}(\mathcal{N}_{1},\delta/2)$, which is the base case $k=1$. Let $j\in\{1,\ldots,k-1\}$ be the smallest integer such that $z^{\prime},g(z^{\prime}),\ldots g^{j-1}(z^{\prime})\in\mathcal{N}(\mathcal{N}_{1},\delta/2)$ but $g^{j}(z^{\prime})\notin\mathcal{N}(\mathcal{N}_{1},\delta/2)$. Since $1\leq j\leq k-1$ we have (c), (d), and (e) above are true for $j$.
@@
so, $g^{j}(\Omega_{U})$ is contained $\delta/2$-neighborhood of $g^{j}(z^{\prime})$. If $g^{j}(z^{\prime})\notin\mathcal{N}(\mathcal{N}_{2},\delta/2),$ then $g^{j}(\Omega_{U})$ is disjoint from $\mathcal{N}_{2}$. But that means the overestimate of $g^{j}(\Omega_{U})$ is disjoint from $\mathcal{N}_{2}$, so the overestimate of the subset $g^{j}(\Omega_{L})$ must also be disjoint from $\mathcal{N}_{2}$. This contradicts Step 2(c)’s conclusion for $j$. Hence, we have (iv): $g^{j}(z^{\prime})\in\mathcal{N}(\mathcal{N}_{2},\delta/2)$ for all $j\in\{1,\ldots,k-1\}$, so $g^{k-1}(z^{\prime})\in\mathcal{N}(\mathcal{N}_{2},\delta/2)$. ∎
@@
Again, first note that if we make it to Step 2, then Step 1(a) did compute the neighborhoods $\mathcal{N}_{2},\mathcal{N}_{1}$, a consequence is that the map $p$ is hyperbolic, thus this need not go into the assumptions.
Since we make it to Step 2, and $N\geq N^{\prime}$, we know (i, ii, iii, iv) of Lemma 3.5 hold. With these useful statements, we consider the loop for $k=1,\ldots,k_{N}$ through Steps 2(b, c, d, e), inductively and show each instance of reporting 0 or $1$ is correct. For this proof, we apply the distortion theorem to $z^{\prime}$ and to points in $\Omega_{L}$ that could be potential points in $J$ inside of $\mathcal{U}^{\prime}$. We use Proposition 2.12 and Lemma 2.13.
@@
(e) test whether $|Dg^{k}(z^{\prime})|_{e}\sigma_{N}c(\gamma)>2\beta^{\prime}\sqrt{2}.$
@@
Suppose to the contrary that there is a point $w^{\prime}\in J\cap\mathcal{U}^{\prime}$. Then by (iii) of Lemma 2.13, since we found $|Dg^{k}(z^{\prime})|_{e}\sigma_{N}c(\gamma)>2\beta^{\prime}\sqrt{2}$, we know there is a reference box of $\mathcal{N}_{2}$ inside the underestimate of $g^{k}(\Omega_{U})$.
@@
On the other hand, if $|Dg^{k}(z^{\prime})|_{e}\sigma_{N}c(\gamma)\leq 2\beta^{\prime}\sqrt{2},$ we pass Step 2(e) and are finished with the loop for this $k$. If $k<k_{N}$, we go back up to Step 2(b) considering $k$ and advance to $k+1$. If $k=k_{N}$, and the algorithm did not halt, we have completed the loop and proceed to Step 2(f). Thus, the reporting of 0 or $1$ for $z^{\prime}$ in Steps 2(c),(d),(e) satisfies the conditions of a $2^{N}$-approximation. ∎
@@
Let $p$ be a polynomial map of $\mathbb{C}$ of degree $d\geq 2$. Beginning to apply Algorithm 3.3, suppose we are examining $z^{\prime}$ an ideal point of denominator $2^{N+2}$ for an $N\geq N^{\prime}$. Assume that Algorithm 3.3 arrives at Step 2(f) and thus reports 0. Then this is consistent with the definition of $2^{N}$-approximation of $J$, i.e., $d(z^{\prime},J)\geq 2^{-N-2}$.
@@
Recall again that if we make it to Step 2, then Step 1(a) did halt and successfully computed the neighborhoods $\mathcal{N}_{2},\mathcal{N}_{1}$, hence $p$ is hyperbolic so this need not go into the assumptions.
@@
As we loop through Steps 2(b, c, d, e), we never report 0 because $w^{\prime}\in J\cap\mathcal{U}^{\prime}$ and by Lemma 3.6, the reporting of 0 was valid in those steps. We would have reported $1$ for some $k$ in Step 2(d) if the underestimate for $g^{k}(\Omega_{U})$ entirely contained a reference box. We show now that $k_{N}$ was defined precisely so that this is true for some $k\in\{1,\ldots,k_{N}\}$ if there is a point in $J\cap\mathcal{U}^{\prime}$.
@@
The previous lemmas combine to establish the following:
@@
The first construction in Step 1 is a constant, independent of $N$, which we do only once at the beginning and don’t need to change for higher $N$.
@@
In Step 2(c) we check if $B^{e}(z^{\prime},|Dg^{k}(z^{\prime})|3\cdot 2^{-N-3}(1+\gamma))$ is disjoint from $\mathcal{N}_{2}$.
In Step 2(d), we perform a similar calculation, checking if $B^{e}(z^{\prime},|Dg^{k}(z^{\prime})|4\cdot 2^{-N-3}(1-\gamma))$ contains any box of $\mathcal{N}_{2}$.
Step 2(e) is a simple check whether $|Dg^{k}(z^{\prime})|_{e}2^{-N-3}c(\gamma)>2\beta^{\prime}\sqrt{2}.$

### v2

@@ 3.2. Poly-time computability of $J$
Step 1(a). First, if $p$ is hyperbolic, we can build a collection of boxes on which an iterate $p^{\nu}$ is expanding. That is, we invoke Lemma 3.1, which succeeds if $p$ is hyperbolic, to compute a positive integer $\nu$, dyadic rationals $\delta>0,L>1$, and the set $\mathcal{N}_{1}$ which is the union of a collection of boxes on which $|Dp^{\nu}(z)|>L$. If $p$ is not hyperbolic, this step will never halt. If this step halts, then the algorithm has just proven that $p$ is hyperbolic (see Lemma 3.9 of [3]).
@@
Step 1(b). Next, we want to invoke Proposition 3.2 to construct our “reference boxes” defining $\mathcal{N}_{2}$. To do so, recall the proposition allows us to input a $b>0$ that is an upper bound for the side-length $\beta^{\prime}$ of the boxes defining $\mathcal{N}_{2}$, though, we never said why such a $b$ might be needed. Now, we provide the calculation defining the $b$ we need.
First set $r=r(\delta):=3\delta/4$ (which is a bit arbitrary but this choice is convenient for calculations) and choose any $\gamma\leq 0.1$ (we see why that $\gamma$ is useful in the proof of (iii) in Lemma 3.5). Then use Equation 2.3 defining $\gamma_{r}(a)$ in terms of $a$ and $r$ to determine $a$ from this $r$ and $\gamma$. Now, set $c=c(\gamma)=\dfrac{(1-3\gamma)(1-\gamma)}{({1+\gamma})}$ (matching Lemma 2.13), and using $\gamma$, $c(\gamma)$, and $a$, set:
(3.1) $b:=\min\left\{\frac{\delta\ c(\gamma)}{16\sqrt{2}(1+\gamma)},a\right\}.$
The reason this value of $b$ is needed is demonstrated in the proof of Theorem 1.1 (more precisely, this $b$ is used at the end of the proof of Lemma 3.5).
With $b$ is determined, we now invoke Proposition 3.2 to compute the union $\mathcal{N}_{2}$ of boxes of computable side-length at most $\beta^{\prime}\in(0,b)$, such that (i) $J\subset\mathcal{N}_{2}$, (ii) each box in the union defining $\mathcal{N}_{2}$ contains a point of $J$, (iii) $|Dg(z)|>L>1$ for every $z\in\mathcal{N}_{2}$, and (iv) $d_{H}(\mathcal{N}_{2},\mathcal{N}_{1})>\delta$ and $\mathcal{N}(\mathcal{N}_{2},\delta)\subset\mathcal{N}_{1}$, where the boxes defining $\mathcal{N}_{1}$ and $\delta>0$ were constructed in Step 1(a) as a set on which the iterate $g=p^{\nu}$ is expanding.
Step 1(c). We explain how to draw the $2^{N}$-picture of $J$. (Why it actually is a $2^{N}$ picture is not explained here, but proven following the algorithm.)
First, recall (at the beginning of the algorithm described in Lemma 3.1) we calculated a large box $R,R]^{2}$ containing $J$. So, we can define $h_{J}(N,z^{\prime})=0$ for all $z^{\prime}\in\mathbb{C}\setminus[-R-1,R+1]^{2}$. Now we consider only $z^{\prime}\in[-R-1,R+1]^{2}$.
In the following, it turns out we need $N\geq N^{\prime}$, where $N^{\prime}$ is the smallest positive integer which satisfies both of the following::
(3.2) $2^{-N^{\prime}-1}<a\ \text{ and }\ 2^{-N^{\prime}}<\delta,$
for the $\delta>0$ determined in Step 1(a) and the $a$ chosen in Step 1(b).
To draw a $2^{N}$-picture of $J$ for some $N<N^{\prime}$ we simply draw the more precise $2^{N^{\prime}}$-approximation.
Now consider any $N\geq N^{\prime}$ (so Equation 3.2 holds for $N$). We perform the following for every ideal point $z^{\prime}\in[-R-1,R+1]^{2}$, a dyadic rational with coordinate denominators in $2^{N+2}$.
Recall that $B(z^{\prime},r)$ denotes the ball (box in our norm) about $z^{\prime}$ of radius $r$. For ease of notation, set $\mathcal{U}^{\prime}:=\overline{B(z^{\prime},2^{-N-2})}$ and $2\mathcal{U}^{\prime}:=\overline{B(z^{\prime},2^{-N-1})}$.
@@
Step 2(a). We first calculate the maximum number of iterates that have to be considered. (We use this in the proof of Lemma 3.7.) Define
@@
where $\lceil\cdot\rceil$ is the ceiling function, and recall from Step 1(b), $\gamma\in(0,0.1]$ and $c=c(\gamma)=(1-3\gamma)(1-\gamma)/({1+\gamma})$, $L$ is the lower bound on expansion of $|Dg|$ in $\mathcal{N}_{1}$ (and $\mathcal{N}_{2}$), and $\beta^{\prime}$ is the side-length of the reference boxes defining $\mathcal{N}_{2}$. It is important that $k_{N}$ is linear in $N$.
@@
We explained in Step 1(a) of Algorithm 3.3 (and in Section 3.1) that if Step 1(a) succeeds, then we have established that $p$ is hyperbolic, and built a union of boxes $\mathcal{N}_{1}$ containing $J$ and on which $p$ is expanding, and a union of smaller boxes $\mathcal{N}_{2}$ containing $J$ such that each box contains a point of $J$, and $\mathcal{N}_{1}\supset\mathcal{N}(\mathcal{N}_{2},\delta)$. (Note if $p$ is not hyperbolic, Step 1(a) never halts so we do not make it to the tests Step 1(d, e, or f) to report anything for any $z^{\prime}$, so this is why we do not need to assume $p$ is hyperbolic for the lemma, that is implied by the assumption the algorithm makes a report in Steps 1(d, e, or f).)
@@
Step 1(e) contained the second check. Since every reference box defining $\mathcal{N}_{2}$ contains a point of $J$, if there is a reference box entirely contained in $2\mathcal{U}^{\prime}$, then there is a point of $J$ in $2\mathcal{U}^{\prime}=\overline{B(z^{\prime},2^{-N-1})},$ so $d(J,z^{\prime})\leq 2^{-N-1}$. So we are allowed to report $1$.
@@
(i): First we show $B^{e}(z^{\prime},r)\subset\mathcal{N}_{1}$, in particular $g$ is conformal on $B^{e}(z^{\prime},r)$, for $r=3\delta/4$. Since Step 1(d) did not triggered a report, we know $\mathcal{N}_{2}\cap\mathcal{U}^{\prime}\neq\emptyset$. Thus $d(z^{\prime},\mathcal{N}_{2})\leq 2^{-N-2}$. Recall that $\mathcal{N}_{1}\supset\mathcal{N}(\mathcal{N}_{2},\delta).$ By Equation 3.2, we have $\delta>2^{-N^{\prime}}\geq 2^{-N}$, thus $2^{-N-2}<\delta/4,$ so $d(z^{\prime},\mathcal{N}_{2})\leq\delta/4,$ hence $B^{e}(z^{\prime},3\delta/4=r)\subset\mathcal{N}_{1}$. Finally, recall that $g$ is conformal on $\mathcal{N}_{1}$ since we have $|Dg(z)|>L>1>0$ there.
(ii): Next we show if $w^{\prime}\in\Omega_{L}$, then $d(w^{\prime},\Omega_{U})\geq\sigma_{N}$, and for any $i\in\mathbb{bN}$, $\frac{1}{2}{\rm diam}(g^{i}(B^{e}(w^{\prime},\sigma_{N}))\geq|Dg^{i}(z^{\prime})|_{e}\sigma_{N}c(\gamma).$ We aim at applying Proposition 2.12 and Lemma 2.13 for $g^{i}=(p^{\nu})^{i},$ and $r,a,\gamma$ as defined in Step 1(b) ($r=(3\delta/4)$, $\gamma\leq 0.1$ and $a<r$ based on $\gamma_{r}(a)$ defined by Equation 2.3), and for $z_{0}=z^{\prime}$, $w_{0}=w^{\prime}$ any point of $\Omega_{L}$, and $\sigma=\sigma_{N}$($2^{-N-3}$ as defined in Step 2(a)), see Figure 1, Right. Since we are in Step 2, we know $\mathcal{U}^{\prime}\cap\mathcal{N}_{2}\neq\emptyset$ so $z^{\prime}\in\mathcal{N}(\mathcal{N}_{2},\delta/2)\subset\mathcal{N}_{1}$ and $\Omega_{L}\subset\Omega_{U}\subset\mathcal{N}_{1}$ as well since $2^{-N-1}<\delta/2$. So since by Equation 3.2, we chose $N^{\prime}$ in Step 1(c) so that $2^{-N^{\prime}-1}<a,$ we know $\Omega_{L}\subset B^{e}(z^{\prime},a)$. Hence $w^{\prime}\in B^{e}(z^{\prime},a)$, and Lemma 2.13 yields (ii).
@@
Observe that if the difference of the radii between the underestimate of $g^{i}(\Omega_{U})$ and the overestimate of $g^{i}(\Omega_{L})$ is at least $\beta^{\prime}\sqrt{2},$ then we know that the reference box about $g^{i}(w^{\prime})\in g^{i}(\Omega_{L})$ is contained entirely inside of the underestimate of $g^{i}(\Omega_{U})$. We show this is true. In fact, this difference in radii is at least:
$|Dg^{i}(z^{\prime})|_{e}4\sigma_{N}(1-\gamma)-|Dg^{i}(z^{\prime})|_{e}3\sigma_{N}(1+\gamma)=|Dg^{i}(z^{\prime})|_{e}\sigma_{N}(1-7\gamma).$
We set $c(\gamma)=(1-3\gamma)(1-\gamma)/(1+\gamma),$ so by an easy calculation, by choice of $\gamma\leq 0.1$, we have $(1-7\gamma)\geq c(\gamma)/2$. (Simplifying this inequality leads to a quadratic which opens down, with two roots, one negative, and the positive root has solution $\gamma=-(8-\sqrt{132}/34)\approx 0.1026213$, so any $\gamma\in(0,0.1]$ satisfies this desired inequality).
Thus, $|Dg^{i}(z^{\prime})|_{e}\sigma_{N}(1-7\gamma)\geq|Dg^{i}(z^{\prime})|_{e}\sigma_{N}c(\gamma)/2.$ Therefore, since by hypothesis this latter quantity is $\beta^{\prime}\sqrt{2}$, we know the the minimum difference between the underestimate of $g^{i}(\Omega_{U})$ to the boundary of the overestimate of $g^{i}(\Omega_{L})$, is greater than the Euclidean diameter of a reference box, and we conclude that the reference box about $g^{i}(w^{\prime})$ is contained entirely inside of the underestimate of $g^{i}(\Omega_{U})$.
@@
Now suppose $k\geq 2$ and we have moved through Steps 2(b,c,d,e) for $k$. Hence, the algorithm has made it through these steps for $1\leq j\leq k-1$ without halting. Thus:
@@
Suppose to the contrary that $g^{k-1}(z^{\prime})\notin\mathcal{N}(\mathcal{N}_{1},\delta/2)$. We know $z^{\prime}\in\mathcal{N}(\mathcal{N}_{1},\delta/2)$, which is the base case $k=1$. Let $j\in\{1,\ldots,k-1\}$ be the smallest integer such that $z^{\prime},g(z^{\prime}),\ldots g^{j-1}(z^{\prime})\in\mathcal{N}(\mathcal{N}_{1},\delta/2)$ but $g^{j}(z^{\prime})\notin\mathcal{N}(\mathcal{N}_{1},\delta/2)$. Since $1\leq j\leq k-1$ we have the above statements (c), (d), and (e) are true for $j$.
@@
and thus $g^{j}(\Omega_{U})$ is contained $\delta/2$-neighborhood of $g^{j}(z^{\prime})$. If $g^{j}(z^{\prime})\notin\mathcal{N}(\mathcal{N}_{2},\delta/2),$ then $g^{j}(\Omega_{U})$ is disjoint from $\mathcal{N}_{2}$. But this means that the overestimate of $g^{j}(\Omega_{U})$ is disjoint from $\mathcal{N}_{2}$, so the overestimate of the subset $g^{j}(\Omega_{L})$ must also be disjoint from $\mathcal{N}_{2}$. This contradicts Step 2(c)’s conclusion for $j$. Hence, we have (iv): $g^{j}(z^{\prime})\in\mathcal{N}(\mathcal{N}_{2},\delta/2)$ for all $j\in\{1,\ldots,k-1\}$, so $g^{k-1}(z^{\prime})\in\mathcal{N}(\mathcal{N}_{2},\delta/2)$. ∎
@@
Again, first note that if we make it to Step 2, then Step 1(a) did compute the neighborhoods $\mathcal{N}_{2},\mathcal{N}_{1}$, a consequence is that the map $p$ is hyperbolic, thus hyperbolicity is not required as an assumption.
Since we make it to Step 2, and $N\geq N^{\prime}$ satisfying Equation 3.2, we know that (i), (ii), (iii), (iv) of Lemma 3.5 hold. With these useful statements, we perform the loop for $k=1,\ldots,k_{N}$ through Steps 2(b, c, d, e), inductively and show each instance of reporting 0 or $1$ is correct. For this proof, we apply the distortion theorem to $z^{\prime}$ and to points in $\Omega_{L}$ that could be potential points in $J$ inside of $\mathcal{U}^{\prime}$. We use Proposition 2.12 and Lemma 2.13.
@@
(e) test whether $|Dg^{k}(z^{\prime})|_{e}\cdot\sigma_{N}\cdot c(\gamma)>2\beta^{\prime}\sqrt{2}.$
@@
Suppose to the contrary that there is a point $w^{\prime}\in J\cap\mathcal{U}^{\prime}$. Then by (iii) of Lemma 2.13, since we found $|Dg^{k}(z^{\prime})|_{e}\cdot\sigma_{N}\cdot c(\gamma)>2\beta^{\prime}\sqrt{2}$, we know there is a reference box of $\mathcal{N}_{2}$ inside the underestimate of $g^{k}(\Omega_{U})$.
@@
On the other hand, if $|Dg^{k}(z^{\prime})|_{e}\cdot\sigma_{N}\cdot c(\gamma)\leq 2\beta^{\prime}\sqrt{2},$ we pass Step 2(e) and are finished with the loop for this $k$. If $k<k_{N}$, we go back up to Step 2(b) considering $k$ and advance to $k+1$. If $k=k_{N}$, and the algorithm did not halt, we have completed the loop and proceed to Step 2(f). Thus, the reporting of 0 or $1$ for $z^{\prime}$ in Steps 2(c),(d),(e) satisfies the conditions of a $2^{N}$-approximation. ∎
@@
Let $p$ be a polynomial map of $\mathbb{C}$ of degree $d\geq 2$. By applying Algorithm 3.3 we examine $z^{\prime}$ an ideal point of denominator $2^{N+2}$ for an $N\geq N^{\prime}$. Assume that Algorithm 3.3 arrives at Step 2(f) and thus reports 0. Then this is consistent with the definition of $2^{N}$-approximation of $J$, i.e., $d(z^{\prime},J)\geq 2^{-N-2}$.
@@
Recall again that if we make it to Step 2, then Step 1(a) did halt and successfully computed the neighborhoods $\mathcal{N}_{2},\mathcal{N}_{1}$, hence $p$ is hyperbolic so this doesn’t need to be included in the assumptions.
@@
As we loop through Steps 2(b, c, d, e), we never report 0 because $w^{\prime}\in J\cap\mathcal{U}^{\prime}$ and by Lemma 3.6, the reporting of 0 was valid in those steps. We would have reported $1$ for some $k$ in Step 2(d) if the underestimate for $g^{k}(\Omega_{U})$ entirely contained a reference box. We now show that $k_{N}$ was defined precisely so that this is true for some $k\in\{1,\ldots,k_{N}\}$ if there is a point in $J\cap\mathcal{U}^{\prime}$.
@@
Combining the previous lemmas establishes the following:
@@
The first computation in Step 1 produces a constant, independent of $N$, as it is carried out only once at the beginning and doesn’t have to be repeated for higher $N$.
@@
In Step 2(c) we check if $B^{e}(z^{\prime},|Dg^{k}(z^{\prime})|\cdot 3\cdot 2^{-N-3}\cdot(1+\gamma))$ is disjoint from $\mathcal{N}_{2}$.
In Step 2(d), we perform a similar calculation, checking if $B^{e}(z^{\prime},|Dg^{k}(z^{\prime})|\cdot 4\cdot 2^{-N-3}\cdot(1-\gamma))$ contains any box of $\mathcal{N}_{2}$.
Step 2(e) is a simple check whether $|Dg^{k}(z^{\prime})|_{e}\cdot 2^{-N-3}\cdot c(\gamma)>2\beta^{\prime}\sqrt{2}.$
@@
4. Concluding Remarks
In this paper, we develop a new approach that re-establishes the poly-time computability of Julia sets for hyperbolic complex polynomials, and allows us to conclude lower semi-computability of the hyperbolicity locus for polynomials of fixed degree. The poly-time computability of hyperbolic Julia sets was originally established by Mark Braverman in 2005. Unlike Braverman’s result, which relies heavily on complex-analytic techniques such as the use of the Poincaré metric and Pick’s theorem, our approach is purely dynamical in nature.
Specifically, we computationally establish uniform expansion in the Euclidean metric for a suitably identified iterate of the polynomial. Among other dynamical tools this step uses the characterization of the Julia set as the expanding chain recurrent set. We then apply several well-known dynamical tools, including invariance properties, quantitative shadowing, and distortion estimates. Since these dynamical techniques are readily available for conformal repellers, our results can be adapted to that setting. Moreover, it is our hope that the methods developed in this paper may serve as a foundation for future computational complexity results concerning more general dynamically defined invariant sets, such as hyperbolic saddle-sets, attractors and higher dimensional Julia sets.